LIST OF PARTICIPANTS

Elvio Accinelli U. de la República, Uruguay

elvio@fing.edu.uy

Grégoire Allaire U. Paris 6, France

allaire@ann.jussieu.fr

Felipe Alvarez U. de Chile, Chile

falvarez@dim.uchile.cl

Jorge Amaya U. de Chile, Chile

jamaya@dim.uchile.cl

Hedy Attouch U. Montpellier II, France

attouch@math.univ-montp2.fr

Didier Aussel U. de Perpignan, France

aussel@univ-perp.fr

Dominique Azé U. de Perpignan, France

aze@univ-perp.fr

Jean-Bernard Baillon U. de Lyon I, France

baillon@desargues.univ-lyon1.fr

Mourad Baïou U. de Chile, Chile

baiou@dim.uchile.cl

Michel Balinski Ecole Polytechnique, France

balinski@poly.polytechnique.fr

Oscar Barrientos Endesa S.A., Chile

obarrien@dim.uchile.cl

Maïtine Bergounioux U. d’Orléans, France

Maitine.Bergounioux@univ-orleans.fr

Silvia Bianchi U. Nacional de Rosario, Argentina

sbianchi@fceia.unr.edu.ar

Jean-Marc Bonnisseau U. de Paris I, France

Jean-Marc.Bonnisseau@univ-paris1.fr

Mireille Bougeard U. de Lyon 1, France

                                                    bougeard@hpvlbi.obspm.fr

Rommel Bustinza

U. Nacional de Ingenieria, Perú

                                                    rommelbp@yahoo.com

Manuel Bustos U. Austral, Chile

                                                    mbustos@valdivia.uca.uach.cl

Giuseppe Buttazzo U. di Pisa, Italia

                                                    buttazzo@dm.unipi.it

Alessandri Canchoa U. Nacional Ingeniería, Perú

                                                    william@fc-uni.edu.pe

Roberto Cominetti U. de Chile, Chile

                                                    rcominet@dim.uchile.cl

Miguel Constantino

U. de Chile, Chile

                                                    mconstan@dii.uchile.cl

Oscar Cornejo U. de Talca, Chile

                                                    ocornejo@maule.utalca.cl

José R. Correa U. de Chile, Chile

                                                    jcorrea@dim.uchile.cl

Rafael Correa U. de Chile, Chile

                                                    rcorreaf@dim.uchile.cl

Jean-Pierre Crouzeix

CUST, U. Blaise Pascal, France

                                                    crouzeix@ucfma.univ-bpclermont.fr

Gonzalo Cueto

U. de la Frontera, Chile

                                                    gcuet001@pinhue.ufro.cl

William Echegaray

U. Nacional Ingeniería, Perú

                                                    william@fc-uni.edu.pe

Ivar Ekeland

U. Paris-Dauphine, France

                                                    Ivar.Ekeland@dauphine.fr

Rafael Epstein U. de Chile, Chile

repstein@dii.uchile.cl

Eduardo Fierro U. Católica del Norte, Chile

efierro@socompa.ucn.cl

Michel Frémond E. Nationale Ponts et Chaussées France

Ubaldo Garcia-Palomares U. Simon Bolivar, Venezuela

garciap@usb.ve

Rolando Garciga Otero IMPA, Brazil

rgarciga@impa.br

Michel Goemans CORE, Belgium

goemans@core.ucl.ac.be

Juan A. Gómez CIMAT, Cuba

jagomez@icmf.inf.cu

Martin Groetschel ZIB-Berlin, Germany

groetschel@zib.de

Hector Guimaray U. Nacional Ingeniería, Perú

william@fc-uni.edu.pe

Elias Gutiérrez U. Nacional Ingeniería, Perú

william@fc-uni.edu.pe

Gonzalo Hernández U. Nacional Andrés Bello, Chile

ghernan@abello.unab.cl

Jean-Baptiste Hiriart-Urruty U. Paul Sabatier, France

jbhu@cict.fr

Alejandro Jofré U. de Chile, Chile

ajofre@dim.uchile.cl

Arturo Labbe U. Católica Valparaiso, Chile

jmena@ucv.cl

Marc Lassonde U. des Antilles et de la Guyane, France

Marc.Lassonde@univ-ag.fr

Juan Carlos Lefimil

U. Católica, Chile

                                                 jlefimil@mat.puc.cl

Bernard Lemaire

U. de Montpellier, France

                                                 lemaire@math.univ-montp2.fr

Christian Malivert U. de Limoges, France

                                                 malivert@unilim.fr

Marcela Miguez U. Católica, Chile

                                                 mmiguez@mat.puc.cl

Boris S. Mordukhovich Wayne State University, USA

                                                 boris@math.wayne.edu

Graciela Nasini

U. Nacional de Rosario, Argentina

                                                 nasini@fceia.unr.edu.ar

Paulo R. Oliveira

UFRJ, Brazil

                                                 poliveir@cos.ufrj.br

Alejandro Omon U. de Chile, Chile

                                                 aomon@dim.uchile.cl

Hector Orellana U. de Chile, Chile

                                                 heorella@dim.uchile.cl

Jaime José Orrillo

IMPA, Brazil

                                                 jjoc@impa.br

Francisco Ortega

CORE-Belgium, U. de Chile

                                                 ortega@core.ucl.ac.be

Carmen Ortíz U. de Chile, Chile

                                                 cortiz@uai.cl

Edwin Palomino

U. Nacional de Ingeniería, Perú

                                                 epalomi@rocketmail.com

Jean-Paul Penot U. de Pau, France

                                                 Jean-Paul.Penot@univ-pau.fr

José-Manuel Pérez-Cerda U. de Chile, Chile

                                                 jmperez@dim.uchile.cl

Pablo Andrés Rey PUC-Rio, Brazil

prey@ele.puc-rio.br

Celso C. Ribeiro PUC-Rio, Brazil

celso@inf.puc-rio.br

Jorge Rivera U. de Chile, Chile

jrivera@dim.uchile.cl

R.T. Rockafellar U. of Washington, USA

rtr@math.washington.edu

Heriberto Román-Flores U. de Tarapacá, Chile

hroman@vitor.faci.uta.cl

Claudia Sagastizábal PUC-Rio, Brazil

sagastiz@ele.puc-rio.br

Dominique Schneider U. d’Amiens, France

msc-math@u-picardie.fr

Geraldo N. Silva IMECC-Unicamp, Brazil

gsilva@nimitz.dcce.ibilce.unesp.br

M. V. Solodov IMPA, Brazil

solodov@impa.br

Wilfredo Sosa Sandoval IMPA, Brazil

sosa@impa.br

Benar Fux Svaiter IMPA, Brazil

benar@impa.br

Michel Théra

U. de Limoges, France

                                                 thera@unilim.fr

Rainer Tichatschke U. Trier, Germany

                                                 tichat@uni-trier.de

Luis Miguel Torres

Escuela Politécnica Nacional, Ecuador

                                                 ltorres@hoy.net

Olga Vasilieva U. del Valle, Colombia

                                                 olgavas@pitagoras.univalle.edu.co

Jorge R. Vera U. Católica, Chile

                                                 jvera@ing.puc.cl

Jean-Philippe Vial U. of Geneva, Switzerland

                                                 jean-philippe.vial@hec.unige.ch

Claire Vilain INRIA/U. de Chile

                                                 vilain@dim.uchile.cl

Michel Volle

U. d’Avignon, France

                                                 michel.volle@univ-avignon.fr

Gabriel Weintraub U. de Chile, Chile

                                                 gweintra@dii.uchile.cl

Roger J-B Wets

U. of California, Davis, USA

                                                 rjbwets@math.ucdavis.edu